OpenStudy (anonymous):

Will medal!!! The addition of hydrochloric acid to a silver nitrate solution precipitates silver chloride according to the reaction: AgNO3(aq)+HCl(aq)→AgCl(s)+HNO3(aq) When you combine 70.0mL of 0.185M AgNO3 with 70.0mL of 0.185M HCl in a coffee-cup calorimeter, the temperature changes from 23.16∘C to 24.25∘C. Calculate ΔHrxn for the reaction as written. Use 1.00 g/mL as the density of the solution and C=4.18J/g⋅∘C as the specific heat capacity.

OpenStudy (jfraser):

seems to me you've been given all the information you need. calculate the \(\Delta\)H using the information given.

OpenStudy (jfraser):

\(\Delta\)H values are given in kj/mol, so you need to find the \(energy\) released by the reaction, using \[Q = m*C*\Delta T\] and solving for Q. Then find the moles of reactant (either the acid or silver nitrate) by using the molarity and volume\[moles = \frac{mole}{L} * L\]and divide #1 by #2

OpenStudy (anonymous):

C = 4.18 delta T = 24.25 - 23.16 = 1.09 mol = 0.070L * 0.185M = 0.01295 m = d*v = 1.00*(0.070+0.070) = 0.14g So, delta H = q/mol HCl = [mC(deltaT)]/mol HCl = (0.14*4.18*1.09)/(0.01295) = 49.3 kJ Thank you so much JFraser! I got it right.